40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1985     Entry time: Fri Sep 11 17:11:15 2009
Author: Sanjit 
Type: Update 
Category: ASS 
Subject: OAF: progress made 

[Jenne & Sanjit]

Good news: We could successfully send filtered output to MC1 @ SUS.

We used 7 channels (different combinations of 3 seismometer and six accelerometer)

We tried some values of \mu (0.001-0.005) & gain on SUS_MC1_POSITION:MCL and C1ASS_TOP_SUS_MC1 (0.1-1).

C1:ASS-TOP_SUS_MC1_INMON is huge (soon goes up to few times 10000), so ~0.1 gains at two places bring it down to a reasonable value.

Bad news: no difference between reference and filtered IOO-MC_L power spectra so far.

Plan of action: figure out the right values of the parameters (\mu, \tau, different gains, and may be some delays), to make some improvement to the spectra.

 ** Rana: there's no reason to adjust any of the MCL gains. We are not supposed to be a part of the adaptive algorithm.

ELOG V3.1.3-